# Kinetic theory of point vortices at order ${1/N}$ and ${1/N^{2}}$

### v1

@@ Abstract
We investigate the long-term relaxation of a distribution of $N$ point vortices in two-dimensional hydrodynamics, in the limit of weak collective amplification. Placing ourselves within the limit of an average axisymmetric distribution, we stress the connections with generic long-range interacting systems, whose relaxation is described within angle-action coordinates. In particular, we emphasise the existence of two regimes of relaxation, depending on whether the system’s profile of mean angular velocity (frequency) is a non-monotonic [resp. monotonic] function of radius, which we refer to as profile (1) [resp. profile (2)]. For profile (1), relaxation occurs through two-body non-local resonant couplings, i.e. $1/N$ effects, as described by the inhomogeneous Landau equation. For profile (2), the impossibility of such two-body resonances submits the system to a “kinetic blocking”. Relaxation is then driven by three-body couplings, i.e. ${1/N^{2}}$ effects, whose associated kinetic equation has only recently been derived. For both regimes, we compare extensively the kinetic predictions with large ensemble of direct $N$-body simulations. In particular, for profile (1), we explore numerically an effect akin to “resonance broadening” close to the extremum of the angular velocity profile. Quantitative description of such subtle nonlinear effects will be the topic of future investigations.
@@ I Introduction
Two-dimensional (${2D}$) geophysical and astrophysical flows have the striking property of organizing spontaneously into coherent structures (large-scale vortices) Chavanis (2002); Bouchet and Venaille (2012). A famous example of this self-organization is Jupiter’s great red spot. Since these coherent structures result from a turbulent mixing, it has sometimes been said that “order emerges from chaos”. To account for this self-organization, Onsager Onsager (1949) considered the statistical mechanics of a system of point vortices of circulations $\gamma_{i}$ (${i\!=\!1,...,N}$) on the plane. He showed the existence of negative temperature states [^1] This was a few years before the discovery of negative temperature states in nuclear spin systems Purcell and Pound (1951). In his visionary paper, Onsager also predicted, in a famous footnote, the quantization of the circulation of point vortices in units of ${h/m}$ in superfluid helium. at which point vortices of the same sign tend to group themselves into “supervortices”. His qualitative arguments were later developed by Joyce and Montgomery Joyce and Montgomery (1973); Montgomery and Joyce (1974), Kida Kida (1975) and Pointin and Lundgren Pointin and Lundgren (1976); Lundgren and Pointin (1977) in a mean field approximation valid for ${N\!\to\!+\infty}$, with ${\gamma\!\sim\!1/N}$. Overall, this leads to the Boltzmann distribution coupled to the Poisson equation. [^2] In unpublished notes Eyink and Sreenivasan (2006), Onsager had previously developed this mean field theory by analogy with his research on electrolytes, which are also described by the Boltzmann–Poisson equations. This equilibrium state (most probable state) can be obtained by maximizing the Boltzmann entropy at fixed circulation, energy, and angular momentum. Later, Miller Miller (1990), as well as Robert and Sommeria Robert and Sommeria (1991), considered the statistical mechanics of more realistic continuous vorticity fields (instead of singular point vortices) described by the ${2D}$ Euler–Poisson equations. In particular, they derived a generalization of the Boltzmann distribution involving a continuum of vorticity levels. For a long time, the distinction between these two statistical theories was not fully understood and appreciated. It was clarified only later by developing the analogy between ${2D}$ vortices and stellar systems Chavanis et al. (1996); Chavanis (2002).
@@
Indeed, self-gravitating systems also self-organize into coherent structures like star clusters, galaxies, or even clusters of galaxies Binney and Tremaine (2008). Stellar systems, which are fundamentally discrete (i.e. made of a collection of $N$ stars in gravitational interaction), undergo two successive types of evolution. In a first regime, one can ignore gravitational encounters between stars and can make a mean field approximation. In this collisionless regime, the locally averaged distribution function of stellar systems is governed by the Vlasov–Poisson equations. Collisionless stellar systems generically experience a process of violent relaxation towards a quasistationary (virialised) state which is a stable stationary solution of the Vlasov–Poisson equations. Lynden-Bell Lynden-Bell (1967) tried to predict this quasistationary state in terms of statistical mechanics but his prediction is hampered by the problem of incomplete relaxation Lynden-Bell (1967); Chavanis (2006). Then, on longer timescales, stellar systems evolve under the effect of gravitational encounters (finite-$N$ effects, granularities, correlations…). During this collisional regime, stellar systems have the tendency to become isothermal but their relaxation towards a true statistical equilibrium state described by the Boltzmann–Poisson equation [^3] Actually, there is no statistical equilibrium state for self-gravitating systems in ${3D}$ because the Boltzmann distribution coupled to the Poisson equation has an infinite mass. Said differently, there is no maximum entropy state at fixed mass and energy. is spoiled by physical effects such as the evaporation of high energy stars and the gravothermal catastrophe (core collapse) Lynden-Bell and Wood (1968).
@@
Similar results apply to ${2D}$ vortices. A gas of point vortices, like a gas of stars, undergoes two successive types of evolution. In a first regime, one can ignore finite-$N$ effects and can make a mean field approximation. In this collisionless limit, the locally averaged vorticity of ${2D}$ point vortices is governed by the ${2D}$ Euler–Poisson equations. Collisionless point vortices generically experience a process of violent relaxation towards a quasistationary state which is a stable stationary solution of the ${2D}$ Euler–Poisson equations. The Miller–Robert–Sommeria Miller (1990); Robert and Sommeria (1991) statistical theory in fluid mechanics, which applies to this regime, is the counterpart of the Lynden-Bell theory Lynden-Bell (1967) in stellar dynamics. Its predictions are also affected by the problem of incomplete relaxation. Then, on longer timescales (collisional regime), ${2D}$ point vortices, just like stellar systems, evolve under the effect of “collisions” towards the statistical equilibrium state described by the Boltzmann–Poisson equation Joyce and Montgomery (1973); Montgomery and Joyce (1974); Kida (1975); Pointin and Lundgren (1976); Lundgren and Pointin (1977).
@@
One can then develop the kinetic theory of ${2D}$ point vortices by analogy with the kinetic theory of stellar systems pioneered by Chandrasekhar Chandrasekhar (1942, 1943a, 1943b, 1949). By making a thermal bath approximation and using the methods of Brownian theory Chandrasekhar (1943c), Chandrasekhar showed that a test star experiences a stochastic process governed by a Fokker–Planck equation involving a diffusion term and a friction term. Similarly, under the same assumptions, a test vortex experiences a stochastic process governed by a Fokker–Planck equation involving a diffusion term and a drift term Chavanis (1998, 2001). As such, the drift of a point vortex is the counterpart of the dynamical friction undergone by a star. These Fokker–Planck equations relax both towards the Boltzmann distribution. In addition, the diffusion and friction (or drift) coefficients are related to each other and to the temperature through some Einstein relation. In particular, in the case of point vortices, the temperature may be negative.
@@
However, this thermal bath approach is oversimplified and does not correctly describe the collisional evolution of the system as a whole. In particular, the Fokker–Planck equation does not conserve the total energy. A better kinetic equation for systems with long-range interactions is the Landau equation Landau (1936) or the Balescu–Lenard equation Balescu (1960); Lenard (1960) which were originally introduced for homogeneous plasmas and applied in a slightly different form to stellar systems by making a local approximation Rosenbluth et al. (1957). These kinetic equations take into account two-body correlations (collisions), so they are valid at order ${1/N}$. In the homogeneous case, they relax towards the Maxwell-Boltzmann distribution on a timescale ${(N/\ln N)\,T_{\mathrm{dyn}}}$, due to logarithmic corrections. These kinetic equations were later extended to describe the axisymmetric evolution of a point vortex gas as a whole Dubin and O’Neil (1988); Dubin (2003); Chavanis (2001); Chavanis and Lemou (2007); Chavanis (2008, 2010, 2012a, 2012b); Fouvry et al. (2016); Chavanis (2023). It turns out that their structure is seemingly different from the original (homogeneous) Landau and Balescu–Lenard equations. Indeed, they involve a resonance condition on the profile of angular velocity ${\Omega(r,t)}$. As long as the profile of angular velocity is non-monotonic, the resonance condition can be satisfied and the system relaxes through two-body effects. However, the relaxation stops when the profile of angular velocity becomes monotonic, even if the system has not reached the statistical equilibrium state. This is an example of “kinetic blocking” Chavanis (2001); Chavanis and Lemou (2007). A similar situation of kinetic blocking had been previously observed for ${1D}$ homogeneous plasmas Eldridge and Feix (1962). In that case, the Balescu–Lenard collision term vanishes identically and the evolution of the system is governed by a collision term of order ${1/N^{2}}$ that accounts for three-body correlations. An explicit expression for this collision term has been obtained recently in the case of homogeneous ${1D}$ systems with long-range interactions, when collective effects are neglected Fouvry et al. (2019, 2020). It always drives the relaxation towards the Boltzmann distribution, showing that the relaxation time scales as ${N^{2}T_{\mathrm{dyn}}}$ in that case.
@@
In reality, stellar systems are spatially inhomogeneous. Their kinetic theory is described by the inhomogeneous Balescu–Lenard equation written with angle-action variables Heyvaerts (2010); Chavanis (2012c). When collective effects are neglected (i.e. for sufficiently hot systems), it reduces to the inhomogeneous Landau equation Chavanis (2013). These equations involve a resonance condition that depends on the orbital frequencies, ${\mathbf{\Omega}(\mathbf{J},t)}$, which themselves are functions of the actions, $\mathbf{J}$. This is similar to the kinetic equation for ${2D}$ point vortices with more general variables. For ${1D}$ inhomogeneous long-range interacting systems with a monotonic frequency profile in which only ${1\!:\!1}$ resonances are permitted, there can also be a kinetic blocking, just like for axisymmetric point vortices. In that case, one has to develop the kinetic theory at the order ${1/N^{2}}$. The expression of the collision term in the absence of collective effects has been obtained in Fouvry (2022). Except for very special interaction potentials Fouvry and Roule (2023), this equation drives the system towards the Boltzmann distribution on a timescale ${N^{2}\,T_{\mathrm{dyn}}}$.
@@
First, we focus on a case with a non-monotonic distribution of orbital frequency (i.e. the vortices’ angular velocity) [profile (1)]. In that case, in the limit where collective effects can be neglected, relaxation can be described with the inhomogeneous Landau equation. We compare quantitatively its prediction with measurements in direct $N$-body simulations, focusing in particular on the dependence of the relaxation rate with the total number of particles and their active fraction (i.e. the fraction of self-consistency in the mean potential). Importantly, we unveil how, close to the extremum of the angular velocity profile, the effective relaxation in the $N$-body simulations is reduced compared to the kinetic prediction. We attribute this effect to “resonance broadening” through nonlinear effects and present a heuristic correction to the Landau equation that closely matches the numerical measurements. Surprisingly, we show that as one delays the impact of nonlinearities, e.g., by increasing the number of particles and/or lowering the active fraction, the resonance broadening seems to saturate and the convergence toward the Landau prediction halts.
Second, we focus on the case of an axisymmetric distribution of point vortices with a monotonic distribution of orbital frequency. In that case, both the ${1/N}$ Landau and ${1/N}$ Balescu–Lenard equations predict a vanishing flux: a “kinetic blocking” (see, e.g., Chavanis, 2001; Chavanis and Lemou, 2007). Placing ourselves once again in the limit of weak collective amplification, we compare quantitatively measurements of the relaxation rate in numerical simulations with the prediction from the recently derived ${1/N^{2}}$ Landau equation Fouvry (2022). We obtain a satisfying agreement between both, recovering in particular the appropriate dependence with respect to the total number of particles and their active fraction.
The paper is organised as follows. In Sec. II, we detail our system. In Sec. III, we investigate long-term dynamics as driven by ${1/N}$ effects [profile (1)], while in Sec. IV, focusing on systems undergoing a kinetic blocking [profile (2)], we consider ${1/N^{2}}$ effects. We conclude in Sec. V. Throughout the main text, technical details are kept to a minimum and deferred to Appendices or to relevant references.
@@ II.1 Hamiltonian
We consider a ${2D}$ system composed of $N$ point vortices of individual circulation ${\gamma_{i}\!=\!\Gamma_{\mathrm{tot}}/N}$, with $\Gamma_{\mathrm{tot}}$ the system’s total active circulation. The dynamics of these vortices is given by the Kirchhoff–Hamilton equations (Newton, 2001)
@@ II.2 Axisymmetric limit
When averaged over independent realisations, the system’s statistics is described via its distribution function (DF), ${F\!=\!F(\mathbf{r},t)}$, normalised to ${\!\int\!\mathrm{d}\mathbf{r}F\!=\!\Gamma_{\mathrm{tot}}}$. We assume that the mean DF, ${F({\bf r},t)\!=\!\langle\sum_{i}\gamma_{i}\delta_{\mathrm{D}}[\mathbf{r}\!-\!\mathbf{r}_{i}(t)]\rangle}$ is axisymmetric so that one has ${F\!=\!F(r,t)}$ with ${r\!=\!|\mathbf{r}|}$. Subsequently, the mean individual Hamiltonian ${H(\mathbf{r},t)\!=\!\int\mathrm{d}\mathbf{r}^{\prime}U(\mathbf{r},\mathbf{r}^{\prime})F(\mathbf{r}^{\prime},t)}$ is ${H\!=\!H(r,t)}$. We can then introduce appropriate angle-action coordinates ${(\theta,J)}$, with $\theta$ the ${2\pi}$-periodic angle and $J$ the associated action, via [^4] These follow from the generating function ${S(y,\theta)\!=\!\tfrac{1}{2}y^{2}\tan\theta}$, to be used in eq. (D.94) of Binney and Tremaine (2008).
@@
Importantly, volume conservation ensures that ${\mathrm{d}\mathbf{r}\!=\!\mathrm{d}\theta\mathrm{d}J}$. With such a choice, it is now clear how an axisymmetric distribution of point vortices can simply be interpreted as one realisation of a long-range interacting inhomogeneous systems (see, e.g., Chavanis, 2012c), characterised by some mean DF, ${F\!=\!F(J,t)}$. Along the same line, from the mean Hamiltonian, ${H\!=\!H(J,t)}$, we define the mean orbital frequency, ${\Omega(J,t)\!=\!\partial H/\partial J}$ [^5] In the original point vortex model (see, e.g., Chavanis, 2012a), the DF ${F(\mathbf{r},t)}$ corresponds to the vorticity ${\omega(\mathbf{r},t)}$, the individual Hamiltonian ${H(\mathbf{r},t)}$ corresponds to the stream function ${\psi(\mathbf{r},t)}$, and the orbital frequency ${\Omega(J,t)}$ corresponds to the angular velocity ${\Omega(r,t)}$ (for axisymmetric flows). For the $2D$ Newtonian interaction potential, one has ${\omega\!=\!-\Delta\psi}$ and ${\Omega(r,t)\!=\!\frac{1}{r^{2}}\int_{0}^{r}\mathrm{d}r^{\prime}\,r^{\prime}\omega(r^{\prime},t)}$. In Fig. 1, we illustrate the typical evolution in configuration space and action space of such axisymmetric distributions.
@@
Figure 1: Left: Illustration in configuration space, ${(x,y)}$, of the dynamics of a few point vortices following the DF from Eq. (6). Right: Same, but in action space ${(\theta,J)}$. In that space, the vortices follow (perturbed) straight-line trajectories.

### v2

# Kinetic theory of two-dimensional point vortices at order 1/N{1/N} and $1/N^{2}$}

@@ Abstract
We investigate the long-term relaxation of a distribution of $N$ point vortices in two-dimensional hydrodynamics. To focus on the regime of weak collective amplification, we embed these point vortices within a static background potential and soften their pairwise interaction on small scales. Placing ourselves within the limit of an average axisymmetric distribution, we stress the connections with generic long-range interacting systems, whose relaxation is described within angle-action coordinates. In particular, we emphasise the existence of two regimes of relaxation, depending on whether the system’s profile of mean angular velocity (frequency) is a non-monotonic [resp. monotonic] function of radius, which we refer to as profile (1) [resp. profile (2)]. For profile (1), relaxation occurs through two-body non-local resonant couplings, i.e. $1/N$ effects, as described by the inhomogeneous Landau equation. For profile (2), the impossibility of such two-body resonances submits the system to a “kinetic blocking”. Relaxation is then driven by three-body couplings, i.e. $1/N^{2}$} effects, whose associated kinetic equation has only recently been derived. For both regimes, we compare extensively the kinetic predictions with large ensemble of direct $N$-body simulations. In particular, for profile (1), we explore numerically an effect akin to “resonance broadening” close to the extremum of the angular velocity profile. Quantitative description of such subtle nonlinear effects will be the topic of future investigations.
@@ I Introduction
Two-dimensional (2D{2D}) geophysical and astrophysical flows have the striking property of organising spontaneously into coherent structures (large-scale vortices) Chavanis (2002); Bouchet and Venaille (2012). A famous example of this self-organization is Jupiter’s great red spot. Since these coherent structures result from a turbulent mixing, it has sometimes been said that “order emerges from chaos”. To account for this self-organization, Onsager Onsager (1949) considered the statistical mechanics of a system of point vortices of circulations $\gamma_{i}$ (i=1,,N{i\!=\!1,...,N}) on the plane. He showed the existence of negative temperature states [^1] This was a few years before the discovery of negative temperature states in nuclear spin systems Purcell and Pound (1951). In his visionary paper, Onsager also predicted, in a famous footnote, the quantization of the circulation of point vortices in units of h/m{h/m} in superfluid helium. at which point vortices of the same sign tend to group themselves into “supervortices”. His qualitative arguments were later developed by Joyce and Montgomery Joyce and Montgomery (1973); Montgomery and Joyce (1974), Kida Kida (1975) and Pointin and Lundgren Pointin and Lundgren (1976); Lundgren and Pointin (1977) in a mean field approximation valid for $N\!\to\!+\infty$}, with $\gamma\!\sim\!1/N$}. Overall, this leads to the Boltzmann distribution coupled to the Poisson equation. [^2] In unpublished notes Eyink and Sreenivasan (2006), Onsager had previously developed this mean field theory by analogy with his research on electrolytes, which are also described by the Boltzmann–Poisson equations. This equilibrium state (most probable state) can be obtained by maximising the Boltzmann entropy at fixed circulation, energy, and angular momentum. Later, Miller Miller (1990), as well as Robert and Sommeria Robert and Sommeria (1991), considered the statistical mechanics of more realistic continuous vorticity fields (instead of singular point vortices) described by the 2D{2D} Euler–Poisson equations. In particular, they derived a generalization of the Boltzmann distribution involving a continuum of vorticity levels. For a long time, the distinction between these two statistical theories was not fully understood and appreciated. It was clarified only later by developing the analogy between 2D{2D} vortices and stellar systems Chavanis et al. (1996); Chavanis (2002).
@@
Indeed, self-gravitating systems also self-organize into coherent structures like star clusters, galaxies, or even clusters of galaxies Binney and Tremaine (2008). Stellar systems, which are fundamentally discrete (i.e. made of a collection of $N$ stars in gravitational interaction), undergo two successive types of evolution. In a first regime, one can ignore gravitational encounters between stars and make a mean field approximation. In this collisionless regime, the locally averaged distribution function of stellar systems is governed by the Vlasov–Poisson equations. Collisionless stellar systems generically experience a process of violent relaxation towards a quasistationary (virialised) state which is a stable stationary solution of the Vlasov–Poisson equations. Lynden-Bell Lynden-Bell (1967) tried to predict this quasistationary state in terms of statistical mechanics but his prediction is spoiled by the problem of incomplete relaxation Lynden-Bell (1967); Chavanis (2006). Then, on longer timescales, stellar systems evolve under the effect of gravitational encounters (finite-$N$ effects, granularities, correlations…). During this collisional regime, stellar systems have the tendency to become isothermal but their relaxation towards a true statistical equilibrium state described by the Boltzmann–Poisson equation [^3] Actually, there is no statistical equilibrium state for self-gravitating systems in 3D{3D} because the Boltzmann distribution coupled to the Poisson equation has an infinite mass. Said differently, there is no maximum entropy state at fixed mass and energy. is hampered by physical effects such as the evaporation of high energy stars and the gravothermal catastrophe (core collapse) Lynden-Bell and Wood (1968).
@@
Similar results apply to 2D{2D} vortices. A gas of point vortices, like a gas of stars, undergoes two successive types of evolution. In a first regime, one can ignore finite-$N$ effects and make a mean field approximation. In this collisionless limit, the locally averaged vorticity of 2D{2D} point vortices is governed by the 2D{2D} Euler–Poisson equations. Collisionless point vortices generically experience a process of violent relaxation towards a quasistationary state which is a stable stationary solution of the 2D{2D} Euler–Poisson equations. The Miller–Robert–Sommeria Miller (1990); Robert and Sommeria (1991) statistical theory in fluid mechanics, which applies to this regime, is the counterpart of the Lynden-Bell theory Lynden-Bell (1967) in stellar dynamics. Its predictions are also affected by the problem of incomplete relaxation. Then, on longer timescales (collisional regime), 2D{2D} point vortices, just like stellar systems, evolve under the effect of “collisions” towards the statistical equilibrium state described by the Boltzmann–Poisson equation Joyce and Montgomery (1973); Montgomery and Joyce (1974); Kida (1975); Pointin and Lundgren (1976); Lundgren and Pointin (1977).
@@
One can then develop the kinetic theory of 2D{2D} point vortices by analogy with the kinetic theory of stellar systems pioneered by Chandrasekhar Chandrasekhar (1942, 1943a, 1943b, 1949). By making a thermal bath approximation and using the methods of Brownian theory Chandrasekhar (1943c), Chandrasekhar showed that a test star experiences a stochastic process governed by a Fokker–Planck equation involving a diffusion term and a friction term. Similarly, under the same assumptions, a test vortex experiences a stochastic process governed by a Fokker–Planck equation involving a diffusion term and a drift term Chavanis (1998, 2001). As such, the drift of a point vortex is the counterpart of the dynamical friction undergone by a star. These Fokker–Planck equations relax both towards the Boltzmann distribution. In addition, the diffusion and friction (or drift) coefficients are related to each other and to the temperature through some Einstein relation. In particular, in the case of point vortices, the temperature may be negative.
@@
However, this thermal bath approach is oversimplified and does not correctly describe the collisional evolution of the system as a whole. In particular, the Fokker–Planck equation does not conserve the total energy. A better kinetic equation for systems with long-range interactions is the Landau equation Landau (1936) or the Balescu–Lenard equation Balescu (1960); Lenard (1960) which were originally introduced for homogeneous plasmas and applied in a slightly different form to stellar systems by making a local approximation Rosenbluth et al. (1957). These kinetic equations take into account two-body correlations (collisions), so they are valid at order 1/N{1/N}. In the homogeneous case, they relax towards the Maxwell-Boltzmann distribution on a timescale $(N/\ln N)\,T_{\mathrm{dyn}}$}, due to logarithmic corrections. These kinetic equations were later extended to describe the axisymmetric evolution of a point vortex gas as a whole Dubin and O’Neil (1988); Dubin (2003); Chavanis (2001); Chavanis and Lemou (2007); Chavanis (2008, 2010, 2012a, 2012b); Fouvry et al. (2016); Chavanis (2023). It turns out that their structure is seemingly different from the original (homogeneous) Landau and Balescu–Lenard equations. Indeed, they involve a resonance condition on the profile of angular velocity $\Omega(r,t)$}. As long as the profile of angular velocity is non-monotonic, the resonance condition can be satisfied and the system relaxes through two-body effects. However, the relaxation stops when the profile of angular velocity becomes monotonic, even if the system has not reached the statistical equilibrium state. This is an example of “kinetic blocking” Chavanis (2001); Chavanis and Lemou (2007). A similar situation of kinetic blocking had been previously observed for 1D{1D} homogeneous plasmas Eldridge and Feix (1962). In that case, the Balescu–Lenard collision term vanishes identically and the evolution of the system is governed by a collision term of order $1/N^{2}$} that accounts for three-body correlations. An explicit expression for this collision term has been obtained recently in the case of homogeneous 1D{1D} systems with long-range interactions, when collective effects are neglected Fouvry et al. (2019, 2020). It always drives the relaxation towards the Boltzmann distribution, showing that the relaxation time scales as $N^{2}T_{\mathrm{dyn}}$} in that case.
@@
In reality, stellar systems are spatially inhomogeneous. Their kinetic theory is described by the inhomogeneous Balescu–Lenard equation written with angle-action variables Heyvaerts (2010); Chavanis (2012c). When collective effects are neglected (i.e. for sufficiently hot systems), it reduces to the inhomogeneous Landau equation Chavanis (2013). These equations involve a resonance condition that depends on the orbital frequencies, $\mathbf{\Omega}(\mathbf{J},t)$}, which themselves are functions of the actions, $\mathbf{J}$. This is similar to the kinetic equation for 2D{2D} point vortices with more general variables. For 1D{1D} inhomogeneous long-range interacting systems with a monotonic frequency profile in which only 1:1{1\!:\!1} resonances are permitted, there can also be a kinetic blocking, just like for axisymmetric point vortices. In that case, one has to develop the kinetic theory at the order $1/N^{2}$}. The expression of the collision term in the absence of collective effects has been obtained in Fouvry (2022). Except for very special interaction potentials Fouvry and Roule (2023), this equation drives the system towards the Boltzmann distribution on a timescale $N^{2}\,T_{\mathrm{dyn}}$}.
@@
First, we focus on a case with a non-monotonic distribution of orbital frequency (i.e. the vortices’ angular velocity) [profile (1)]. To place ourselves in the limit where collective effects can be neglected, we embed the point vortices within a static background potential and soften the pairwise interaction on small scales. In that case, relaxation can be described with the inhomogeneous Landau equation. We compare quantitatively its prediction with measurements in direct $N$-body simulations, focusing in particular on the dependence of the relaxation rate with the total number of particles and their active fraction (i.e. the fraction of self-consistency in the mean potential). Importantly, we unveil how, close to the extremum of the angular velocity profile, the effective relaxation in the $N$-body simulations is reduced compared to the kinetic prediction. We attribute this effect to “resonance broadening” through nonlinear effects and present a heuristic correction to the Landau equation that closely matches the numerical measurements. Surprisingly, we show that as one delays the impact of nonlinearities, e.g., by increasing the number of particles and/or lowering the active fraction, the resonance broadening seems to saturate and the convergence toward the Landau prediction halts.
Second, we focus on the case of an axisymmetric distribution of point vortices with a monotonic distribution of orbital frequency. In that case, both the 1/N{1/N} Landau and 1/N{1/N} Balescu–Lenard equations predict a vanishing flux: a “kinetic blocking” (see, e.g., Chavanis, 2001; Chavanis and Lemou, 2007). Once again, we place ourselves in the limit of weak collective amplification by embedding the point vortices within a static background potential and by softening their pairwise interaction on small scales. In that dynamically hot regime, we compare quantitatively measurements of the relaxation rate in numerical simulations with the prediction from the recently derived $1/N^{2}$} Landau equation Fouvry (2022). We obtain a satisfying agreement between both, recovering in particular the appropriate dependence with respect to the total number of particles and their active fraction.
The paper is organised as follows. In Section II, we detail our system. In Section III, we investigate long-term dynamics as driven by 1/N{1/N} effects [profile (1)], while in Section IV, focusing on systems undergoing a kinetic blocking [profile (2)], we consider $1/N^{2}$} effects. We conclude in Section V. Throughout the main text, technical details are kept to a minimum and deferred to Appendices or to relevant references.
@@ II.1 Hamiltonian
We consider a 2D{2D} system composed of $N$ point vortices of individual circulation $\gamma_{i}\!=\!\Gamma_{\mathrm{tot}}/N$}, with $\Gamma_{\mathrm{tot}}$ the system’s total active circulation. The dynamics of these vortices is given by the Kirchhoff–Hamilton equations (Newton, 2001)
@@ II.2 Axisymmetric limit
When averaged over independent realisations, the system’s statistics is described via its distribution function (DF), $F\!=\!F(\mathbf{r},t)$}, normalised to $\int\!\mathrm{d}\mathbf{r}F\!=\!\Gamma_{\mathrm{tot}}$}. We assume that the mean DF, $F({\bf r},t)\!=\!\langle\sum_{i}\gamma_{i}\delta_{\mathrm{D}}[\mathbf{r}\!-\!\mathbf{r}_{i}(t)]\rangle$} is axisymmetric so that one has F=F(r,t){F\!=\!F(r,t)} with $r\!=\!|\mathbf{r}|$}. Subsequently, the mean individual Hamiltonian $H(\mathbf{r},t)\!=\!\int\mathrm{d}\mathbf{r}^{\prime}U(\mathbf{r},\mathbf{r}^{\prime})F(\mathbf{r}^{\prime},t)$} is H=H(r,t){H\!=\!H(r,t)}. We can then introduce appropriate angle-action coordinates $(\theta,J)$}, with $\theta$ the $2\pi$}-periodic angle and $J$ the associated action, via [^4] These follow from the generating function $S(y,\theta)\!=\!\tfrac{1}{2}y^{2}\tan\theta$}, to be used in eq. (D.94) of Binney and Tremaine (2008).
@@
Importantly, volume conservation ensures that $\mathrm{d}\mathbf{r}\!=\!\mathrm{d}\theta\mathrm{d}J$}. With such a choice, it is now clear how an axisymmetric distribution of point vortices can simply be interpreted as one example of a long-range interacting inhomogeneous systems (see, e.g., Chavanis, 2012c), characterised by some mean DF, F=F(J,t){F\!=\!F(J,t)}. Along the same line, from the mean Hamiltonian, H=H(J,t){H\!=\!H(J,t)}, we define the mean orbital frequency, $\Omega(J,t)\!=\!\partial H/\partial J$}. [^5] In the original point vortex model (see, e.g., Chavanis, 2012a), the DF $F(\mathbf{r},t)$} corresponds to the vorticity $\omega(\mathbf{r},t)$}, the individual Hamiltonian $H(\mathbf{r},t)$} corresponds to the stream function $\psi(\mathbf{r},t)$}, and the orbital frequency $\Omega(J,t)$} corresponds to the angular velocity $\Omega(r,t)$} (for axisymmetric flows). For the $2D$ Newtonian interaction potential, one has $\omega\!=\!-\Delta\psi$} and $\Omega(r,t)\!=\!\frac{1}{r^{2}}\int_{0}^{r}\mathrm{d}r^{\prime}\,r^{\prime}\omega(r^{\prime},t)$}. In Fig. 1, we illustrate the typical evolution in configuration space and action space of such axisymmetric distributions.
@@
Figure 1: Left: Illustration in configuration space, (x,y){(x,y)}, of the dynamics of a few point vortices following the DF from Eq. (6). Right: Same, but in action space $(\theta,J)$}. In that space, the vortices follow (perturbed) straight-line trajectories.